# FOF-X: Towards Real-time Detailed Human Reconstruction from a Single Image

### v3

@@ IV-E Ablation Study
IV-E1 Number of Terms to Use
In the original FOF, we use $2N+1=31$ terms of the Fourier series to represent the occupancy field, which has been shown to be sufficient for maintaining geometric accuracy. However, we observe that the absence of high-frequency components can lead to visual artifacts. Therefore, we use $N=128$ terms of the cosine series in FOF-X. We illustrate the impact of varying NN on the accuracy of the approximation in Fig. 10. The metrics for different values of NN are also presented in Table III. When $N\leq 16$, noticeable artifacts appear on the meshes. Noting the geometric details in the ear region, although using $N=32$ preserves most geometric information, we prefer $N=128$ in FOF-X to ensure higher visual quality and eliminate potential artifacts.
IV-E2 Sampling Scalability
The frequency-domain resolution of FOF can be adaptively inferred according to the time needs of the system, without the need for retraining. We explored the effect of resolution (number of components) on the results taking $N=128$. Fig. 11 illustrates the effect of changing resolution on the accuracy of the results. Metrics of different NN values are also given in Table IV.
IV-E3 Noise Robustness
Neural networks often struggle to produce precise results, and their outputs can be considered as samples from a certain distribution, leading to inherent noise in the results. The geometric representation used must therefore be robust to noise and small perturbations. Therefore, representations that are sensitive to noise are not suitable for machine learning. To verify the noise robustness of FOF, we manually add relative Gaussian noise with different variances and show the quantitative results in Fig. 12 and Table V. Based on these experiments, we can conclude that the FOF is very robust to noise and can keep a general shape even with large noise. Thus, it is reasonable and feasible to use neural networks to estimate FOFs.
@@ IV-E3 Noise Robustness
Figure 10: Results of different number NN of coefficients used in FOF-X.
@@
IV-E4 Dual-sided Normal Maps
IV-E5 Improved Inter-Conversion Algorithm
Fig. 2 shows the effectiveness and robustness of our improved inter-conversion algorithm. As can be seen, our parallelized mesh-to-FOF algorithm demonstrates strong robustness. The quantitative results for two mesh-to-FOF algorithms are also given in Table VI. With the automaton-based discontinuity matcher, it effectively avoids the floating masses caused by duplicate or missing discontinuities. This algorithm is integrated into the real-time pipeline for its excellent parallelizability. Table VII shows the quantitative results of the FOF-to-mesh methods. To better represent the visual effect of the mesh, we render the mesh as a normal diagram and introduce visual metrics [^2] SSIM (structural similarity), PSNR (peak signal-to-noise ratio) and LPIPS (learned perceptual image patch similarity) measure structural preservation, pixel-level accuracy and human perceptual consistency, respectively. to evaluate results. Our improved FOF-to-mesh algorithm avoids the stair-step artifacts caused by view bias effectively. Although the Laplacian coordinate constraint solution is a little time-consuming, we can use Laplacian smoothing in the real-time pipeline as an alternative and apply our full FOF-to-mesh algorithm during the mesh export stage.
TABLE V: Geometric Metrics for Different Noise Level
Noise level 0% 5% 10% 15% 20% 25% 30%
P2S 0.027 0.030 0.035 0.041 0.048 0.075 0.224
Chamfer 0.030 0.032 0.037 0.043 0.049 0.065 0.141
Figure 12: Results showing the effects of different noise levels applied to FOF-X, demonstrating FOF-X’s robustness to noise.
IV-F The Real-time Pipeline
To perform real-time monocular human reconstruction, we design a four-stage pipeline. In the first stage, we get an image from the video stream with OpenCV [52], use RVM [53] to get a human mask to remove the background, and predict a SMPL model as prior with 4D-humans[54]. RVM is originally a matting method, and we use a threshold of 0.5 to obtain a binary mask. In the second stage, a CUDA-implemented rasterizer is used to render dual-sided normal maps and a FOF representation for the predicted SMPL. The proposed mesh-to-FOF conversion algorithm is integrated into the rasterizer. In the third stage, the pre-trained dual-sided normal estimator and our reconstruction network are encapsulated into a single neural network module, which is used to reconstruct the FOF. In the final stage, the FOF is resized to a proper resolution ($256\times 256$ in our implementation) and transformed into a mesh. All four stages are implemented in PyTorch and accelerated with TensorRT, running on a single RTX-4090 GPU. Our pipeline works at 30+ FPS, which is limited by the camera (Logic C920 Pro). With a recorded video, it can achieve a frame rate of about 50 FPS. Fig. 13 shows the real-time reconstruction system (all human subjects provided informed consent for data collection). Please check the demo video for more results.
TABLE VI: Geometric Metrics for Different Mesh-to-FOF algorithms
@@ IV-F The Real-time Pipeline
TABLE VII: Visual Metrics for Different FOF-to-mesh algorithms
@@
Figure 13: Our system applied to perform real-time 3D human reconstruction from a single monocular RGB input. Please see the supplement video for the live demonstration.
@@ V Conclusion and discussion
Conclusion. In this paper, we first present the Fourier Occupancy Field (FOF) representation for monocular real-time 3D human reconstruction and further extend it to the FOF-X framework for more detailed results. FOF is an efficient and flexible 3D representation defined as a 2D vector field orthogonal to the $z$-axis. When extending the FOF representation to the FOF-X reconstruction framework, we introduce additional improvements to the conversion algorithm and the reconstruction algorithm. For the conversion algorithm, an automaton-based discontinuity matcher is designed to convert a triangle mesh to a FOF efficiently and robustly. Thanks to this, the SMPL mesh prior can be added into the real-time pipeline, further enhancing the performance of the system. Besides, the FOF to mesh conversion is also improved with the Laplacian coordinate constraint, eliminating the stair-step artifacts. For the reconstruction framework, we predict dual-sided normal maps from input RGB images first and then use dual-sided normal maps instead of RGB images as the input of the main network, which greatly mitigates the performance degradation caused by the domain gap between the training images and the real images. The efficacy of FOF and FOF-X is validated on 3D scanned and in-the-wild datasets, where our approaches effectively recover detailed human geometry with better quality than previous solutions.
Limitations and Future Work. The FOF representation is based on the Fourier series expansion of the square-wave-like function. When applying FOF to thin objects, the spectrum of the function will contain many high-frequency components. In such cases, we need more terms in the Fourier series to approximate the function. The number of terms we need is roughly inversely proportional to the thickness of the object. In our experiments, we found that FOF fails to reconstruct objects that are too thin, like single layer mesh. Although wavelet transformation and short-time Fourier transformation seem to be a promising option, the implementation and computational efficiency still need more exploration. In addition, the FOF representation shows significant potential for broader applications. Its mathematical formulation and efficient parameterization could be extended to general object representation and even scene-level modeling through appropriate architectural modifications. These promising directions will be explored in future work to enhance the framework’s capabilities and applicability.
VI Acknowledgements
This work was partially supported by the National Key R&D Program of China (2023YFC3082100) and the National Natural Science Foundation of China (62171317, 62125107, and 62231018).

### v4

@@ IV-C Ablation Study
IV-C1 Number of Terms to Use
In the original FOF, we use $2N+1=31$ terms of the Fourier series to represent the occupancy field, which has been shown to be sufficient for maintaining geometric accuracy. However, we observe that the absence of high-frequency components can lead to visual artifacts. Therefore, we use $N=128$ terms of the cosine series in FOF-X. We illustrate the impact of varying $N$ on the accuracy of the approximation in Fig. 8. The metrics for different values of $N$ are also presented in Table III. When $N\leq 16$, noticeable artifacts appear on the meshes. Noting the geometric details in the ear region, although using $N=32$ preserves most geometric information, we prefer $N=128$ in FOF-X to ensure higher visual quality and eliminate potential artifacts.
IV-C2 Sampling Scalability
The frequency-domain resolution of FOF can be adaptively inferred according to the time needs of the system, without the need for retraining. We explored the effect of resolution (number of components) on the results taking $N=128$. Fig. 9 illustrates the effect of changing resolution on the accuracy of the results. Metrics of different $N$ values are also given in Table IV.
@@ IV-C2 Sampling Scalability
Figure 8: Results of different number $N$ of coefficients used in FOF-X.
@@
TABLE V: Geometric Metrics for Different Mesh-to-FOF algorithms
@@
TABLE VI: Visual Metrics for Different FOF-to-mesh algorithms
@@
IV-C3 Dual-sided Normal Maps
IV-C4 Improved Inter-Conversion Algorithm
Fig. 2 shows the effectiveness and robustness of our improved inter-conversion algorithm. As can be seen, our parallelized mesh-to-FOF algorithm demonstrates strong robustness. The quantitative results for two mesh-to-FOF algorithms are also given in Table V. With the automaton-based discontinuity matcher, it effectively avoids the floating masses caused by duplicate or missing discontinuities. This algorithm is integrated into the real-time pipeline for its excellent parallelizability. Table VI shows the quantitative results of the FOF-to-mesh methods. To better represent the visual effect of the mesh, we render the mesh as a normal diagram and introduce visual metrics [^2] SSIM (structural similarity), PSNR (peak signal-to-noise ratio) and LPIPS (learned perceptual image patch similarity) measure structural preservation, pixel-level accuracy and human perceptual consistency, respectively. to evaluate results. Our improved FOF-to-mesh algorithm avoids the stair-step artifacts caused by view bias effectively. Although the Laplacian coordinate constraint solution is a little time-consuming, we can use Laplacian smoothing in the real-time pipeline as an alternative and apply our full FOF-to-mesh algorithm during the mesh export stage.
@@ V Conclusion and Discussion
In this paper, We presented the Fourier Occupancy Field (FOF), an efficient 3D representation for monocular real-time human reconstruction, and extended it to the FOF-X framework for higher-quality results. FOF encodes the occupancy field into a compact multi-channel 2D form, enabling simple image-to-image estimation and seamless conversion with meshes. Building on this, we proposed the FOF-X framework with enhanced conversion and reconstruction algorithms. Specifically, an automaton-based discontinuity matcher enables robust mesh-to-FOF conversion and incorporation of SMPL priors into real-time pipelines, while Laplacian coordinate constraints eliminate stair-step artifacts in FOF-to-mesh conversion. For reconstruction, we employ dual-sided normal maps as input, effectively reducing domain gaps between training and real images. Experiments on scanned and in-the-wild datasets confirm that FOF and FOF-X achieve high-quality and detailed human reconstruction, surpassing prior approaches.
The FOF representation is based on the Fourier series expansion of the square-wave-like function. When applying FOF to thin objects, the spectrum of the function will contain many high-frequency components. In such cases, we need more terms in the Fourier series to approximate the function. The number of terms we need is roughly inversely proportional to the thickness of the object. In our experiments, we found that FOF fails to reconstruct objects that are too thin, like single layer mesh. Although wavelet transformation and short-time Fourier transformation seem to be a promising option, the implementation and computational efficiency still need more exploration. In addition, the FOF representation shows significant potential for broader applications. Its mathematical formulation and efficient parameterization could be extended to general object representation and even scene-level modeling through appropriate architectural modifications. These promising directions will be explored in future work to enhance the framework’s capabilities and applicability. We leave this issue for future work to further improve our framework.
Although a previous demo paper of FOF [demo] has demonstrated a system with a perspective camera setting, it encountered challenges in handling changes in the camera parameters. Integrating the camera setting of a true perspective camera with our method for more accurate mesh recovery remains a topic for future research. Moreover, we have noticed that the combination of various 3D representations and generative models has produced a bunch of amazing works. Given the good nature of FOF, combining FOF with large models to sense the general 3D world is also a promising direction.